# iKap: Kinematics-aware Planning with Imperative Learning

### v1

@@ Abstract
Trajectory planning in robotics aims to generate collision-free pose sequences that can be reliably executed. Recently, vision-to-planning systems have garnered increasing attention for their efficiency and ability to interpret and adapt to surrounding environments. However, traditional modular systems suffer from increased latency and error propagation, while purely data-driven approaches often overlook the robot’s kinematic constraints. This oversight leads to discrepancies between planned trajectories and those that are executable. To address these challenges, we propose iKap, a novel vision-to-planning system that integrates the robot’s kinematic model directly into the learning pipeline. iKap employs a self-supervised learning approach and incorporates the state transition model within a differentiable bi-level optimization framework. This integration ensures the network learns collision-free waypoints while satisfying kinematic constraints, enabling gradient back-propagation for end-to-end training. Our experimental results demonstrate that iKap achieves higher success rates and reduced latency compared to the state-of-the-art methods. Besides the complete system, iKap offers a visual-to-planning network that seamlessly integrates kinematics into various controllers, providing a robust solution for robots navigating complex and dynamic environments.
@@ I Introduction
Path planning is a fundamental task in robotics, involving the determination of a collision-free trajectory that connects the robot’s current position to the goal [1]. In dynamic and complex environments, robots must accurately perceive their surroundings and correlate environmental data with the planned path to ensure successful task execution [2]. Recently, with the development of perceptual systems, direct vision-to-planning systems have garnered increasing attention due to their potential efficiency and the ability to interpret and adapt to surrounding environments [3].
@@
Data-driven end-to-end approaches, particularly those utilizing deep learning, have obtained significant advancement [3]. Such methods directly map sensor inputs to control outputs, eliminating the error propagation and latency inherent in modular systems [6, 7, 8, 3]. However, purely data-driven methods often act as black boxes, require vast amounts of data, have low sample efficiency, and exhibit a significant sim-to-real gap. In addition, these approaches lack interpretability, making it challenging to incorporate explicit physical constraints or prior knowledge [9]. Although recent works such as iPlanner [10] have explored combining data-driven methods with trajectory optimization, they neglect the robot’s kinematics. This absence of physical constraints can result in generated trajectories that are difficult for low-level controllers to execute, leading to issues such as deadlocks, collisions, or suboptimal performance [11]. As illustrated in Figure 1, for a robot with kinematic constraints like a minimum turning radius, the planned trajectory (blue curve) may not be executable by the robot.
To address the issue of previous end-to-end systems often generating infeasible paths, we argue that kinematic constraints should be incorporated into vision-to-planning networks while retaining the end-to-end learning architecture to avoid execution errors.
To this end, we propose an Imperative learning-based Kinematics-aware Planning (iKap) system that integrates the robot’s kinematic model directly into the learning pipeline. This system supervises the network to learn kinematically feasible trajectories through self-supervised bi-level optimization (BLO). The learning module of iKap is highly flexible and can be easily integrated with various controllers to plan executable trajectories. Our main contributions include:
  • We integrate the robot’s kinematic model into the training pipeline through a differentiable model predictive control (MPC) using a self-supervised bi-level optimization framework via imperative training. This enables the network to learn the kinematically feasible trajectories.
@@ II-A Data-driven Planner
Our research builds upon iPlanner [10], an end-to-end local planner that formulates trajectory planning as a bi-level optimization problem. iPlanner guides neural networks to plan collision-free paths using an imperative loss based on the Euclidean Signed Distance Field (ESDF). However, it assumes that robots can perfectly execute trajectories optimized purely from geometric constraints, neglecting kinematic feasibility and thus limiting real-world performance, especially for agile robots. Additionally, iPlanner simplifies the bi-level optimization to a single-level problem via a closed-form solution for the lower-level problem, enhancing computational efficiency but reducing generalizability to other constraints. ViPlanner [12] extends iPlanner by incorporating semantic information to construct cost maps for semantic-based local planning, yet it does not address the limitations regarding kinematic feasibility and generalizability.
@@
Differentiable simulation has recently gained increasing attention for its ability to accelerate model learning by leveraging autodiff tools to compute the gradient propagation of system dynamics. Previous studies have demonstrated that integrating differentiable simulation into reinforcement learning systems significantly improves sample efficiency, enabling applications in vision-based control tasks [17]. Furthermore, recent work has integrated a simple point-mass model with a deep learning pipeline and achieved visual-based navigation tasks by introducing temporal gradient decay to mitigate gradient explosion [18].
@@ II-B Differentiable Optimization and Imperative Learning
Traditional robotics algorithms often rely on optimization techniques, which can establish implicit relationships between input and output data. Integrating optimization-based algorithms into end-to-end systems can narrow the search space for neural networks, reduce the number of parameters, and improve computational efficiency [16].
In this context, the differentiable optimization solver is critical, ensuring correct gradient propagation during the learning process. Amos et al. proposed implicit differentiation to manage gradient backpropagation for argmin problems, integrating this process as a layer in neural networks. In their follow-up work, they introduced differentiable MPC, which plays a significant role in bridging optimization with deep learning frameworks [19, 20].
In addition, incorporating differentiable optimization into end-to-end training frameworks also preserves the interpretability of classical methods. This integration offers a promising avenue for developing more sophisticated and interpretable robotic systems. Building on these concepts, imperative learning has been proposed. It incorporates optimization-based algorithms as constraints, adapting to the robot’s state and facilitating self-supervised learning [21]. Ultimately, the network training problem can be formulated as a bi-level optimization. However, a critical challenge is the gradient backpropagation from the upper-level to the lower-level optimization, which was bypassed by many previous works either using closed-form solutions [10], gradient assumptions at convergence [22, 23], or discrete approximations [24, 25]. Our problem relies heavily on accurate gradient due to the necessity of physical constraint in the BLO and strong coupling between the two level optimization. We address this issue by employing a differentiable MPC [26, 27], demonstrating that imperative learning can successfully operate in scenarios with constraint optimization. This confirms the versatility and robustness of this architecture in complex robotic applications.
@@ III-A Problem Formulation and System Overview
Figure 2: The training pipeline for the planner is based on bi-level optimization. In this framework, the networks, the optimized object, generate waypoints based on the given depth perception and goals. Then, the low-level trajectory optimization module tracks a kinematics-feasible trajectory. The embedded kinematics and gradients from the lower-level module are then utilized to supervise and train the networks.
@@
As illustrated in Fig. 2, our iKap consists of three key parts. First, a ResNet[28] front-end takes the depth observation $\mathcal{D}_{t}$ and encodes it into observation embedding. This embedding, combined with the target $\mathbf{p}_{t}^{g}$, is fed into the second planning network to predict a sparse key point trajectory $\boldsymbol{\mu}_{1:k}$ and collision probability $pr_{t}$. Third, a differentiable MPC tracks a reference $\boldsymbol{\tau}:=\mathbf{x}_{1:T+1}\in\mathcal{X}^{T+1}$ interpolated from $\boldsymbol{\mu}_{1:k}$ with time horizon $T$ and outputs the optimized trajectory $\boldsymbol{\tau}^{*}_{1:T+1}$ and action $\mathbf{u}^{*}_{1:T}$ while ensuring the kinematic feasibility $\mathbf{x}_{t+1}=\mathbf{F}(\mathbf{x}_{t},\mathbf{u}_{t})$. Here, $\mathbf{F}$ is the robot kinematics, $\mathbf{x}_{t}$ and $\mathbf{u}_{t}$ are the robot’s states and control input in configuration space $\mathcal{X}$ and action space $\mathcal{A}$, respectively. Finally, a well-designed traversability costs $\mathcal{U}$ is evaluated and backpropagated through both the differentiable MPC and the perception network to update the network parameters $\theta$. This process results in a bi-level optimization (BLO) problem (1), where perception and planning network is jointly optimized with the differentiable MPC.
@@
[…]
@@ III-C Differentiable Model-Predictive Control
[…]
@@
where $\tilde{\mathbf{x}}_{t}=\mathbf{x}_{t}-\mathbf{x}_{t}^{\text{ref}},t=0,\cdots,T$ is the error state between the current state $\mathbf{x}_{t}$ and reference trajectory $\mathbf{x}^{\text{ref}}$; $\tilde{\mathbf{x}}^{\top}_{t}\mathbf{Q}_{t}\tilde{\mathbf{x}}_{t}+\mathbf{u}_{t}^{\top}\mathbf{R}_{t}\mathbf{u}_{t}$ is the stage cost balancing the state and control effort; $\tilde{\mathbf{x}}^{\top}_{T}\mathbf{Q}_{T}\tilde{\mathbf{x}}_{T}$ is the terminal cost; $\mathbf{Q}_{t}$, $\mathbf{R}_{t}$,$\mathbf{Q}_{T}$ are the weight matrices. Specifically, the reference trajectory is obtained by interpolating the waypoints predicted by the planning network. The classic Dubins car (5) can be used to describe kinematic model of a legged robot.
@@
Since $\boldsymbol{\tau}^{*}$ is a function of the reference path $\boldsymbol{\tau}$, which is a function of the predicted key point path $\boldsymbol{\mu}_{1:k}$ or $f_{\theta}$, and will be passed to the upper level optimization for the final end-to-end training, the critical part is to calculate the gradient $\frac{\partial\boldsymbol{\tau}^{*}}{\partial\boldsymbol{\tau}}$. The traditional unrolling approach maintain the computational graph throughout the entire iteration process [21], which poses significant computation burden when dealing with complex problem. It may also run into divergent or vanishing. Instead, we employ the implicit function differentiation theorem and leveraging the KKT condition of (2) at the optimal point to compute the gradients of the parameters [20]. Thus, there is no need for explicit unrolling of the entire iteration process. In practice, we use the differentiable MPC module in our robot learning library PyPose [29, 26], which solves the MPC by the iterative Linear Quadratic Regulator (iLQR) algorithm [27]. PyPose MPC adopts a more general and problem-agnostic automatic differentiation (AD)-based solution that introduces just one additional optimization iteration at the optimal point in the forward pass. This allows automatic gradient calculation during the backward pass without the need to differentiate through the entire unrolled chain or compute problem-specific analytical derivatives.
@@ III-D Upper-level Cost Design and Optimization
The environmental cost $\mathcal{C}^{\mathcal{E}}$ in (8) represents the interaction cost between the environment and the trajectory during planning. The shortest distance between the robot and the environment, i.e., ESDF map, is used to ensure safety.
@@
$\mathcal{C}^{\mathcal{T}}=$ $\gamma_{1}\cdot\log(\|\boldsymbol{\mu}_{k}-\mathbf{p}_{t}^{g}\|_{2}+1.0)$ (9)
@@
The first term encourages the planner to generate a trajectory close to the goal. The second term measures the deviation from a straight, uniformly distributed path, ensuring that the trajectory is evenly spaced. This design also helps smooth out disorganized paths, preventing the learning process from becoming trapped in local minima. The final term focuses on the relationship between the reference and kinematics-feasible trajectory. Minimizing it allows the network to learn the distribution of kinematically feasible trajectories.
To solve the BLO in (1), the gradient of the upper-level cost must be propagated back through the lower-level system by applying the chain rule, and the network parameter is updated using gradient decent:
@@ IV-A Implementation, Platforms, and Baselines
Figure 3: The real-world experiment pipeline. The experiments are conducted on the Unitree Go2 robot dog. We use AirVO to estimate the robot’s odometry, command the goal waypoint to the robot, and inputs this information into iKap along with depth images captured by the RealSense D435i camera.
@@
    iKap+PID: This setting is at test time. We use a PID controller to track iKap planned reference trajecotry.
@@
    iKap+MPC: This setting is at test time. We use an MPC controller to track iKap planned reference trajecotry.
@@ IV-B3 Navigation Test
We next evaluate the planners’ performance in a navigation task. To this end, we calculate the failure rate of a planner, where a failed attempt is defined as any instance where a collision with the environment occurred, the robot became deadlocked during planning, or the trajectory could not be executed by the controller. Specifically, we collected 100 pairs of start poses and goal positions and used both iPlanner and iKap to generate local trajectories and navigate the robot to its destination. As shown in Table II, iKap demonstrates superior performance in three environments. It can be seen that failure rate of iKap is slightly higher than in the “Garage” environment, this is because the robot often faces obstructed vision due to the presence of numerous walls and blind spots, while iKap is sensitive to that kinematic infeasible region.
@@
iPlanner 70.59 73.91 87.10 79.31
iKap 82.35 69.57 93.54 86.20

### v3

@@ Abstract
Trajectory planning in robotics aims to generate collision-free pose sequences that can be reliably executed. Recently, vision-to-planning systems have gained increasing attention for their efficiency and ability to interpret and adapt to surrounding environments. However, traditional modular systems suffer from increased latency and error propagation, while purely data-driven approaches often overlook the robot’s kinematic constraints. This oversight leads to discrepancies between planned trajectories and those that are executable. To address these challenges, we propose iKap, a novel vision-to-planning system that integrates the robot’s kinematic model directly into the learning pipeline. iKap employs a self-supervised learning approach and incorporates the state transition model within a differentiable bi-level optimization framework. This integration ensures the network learns collision-free waypoints while satisfying kinematic constraints, enabling gradient back-propagation for end-to-end training. Our experimental results demonstrate that iKap achieves higher success rates and reduced latency compared to the state-of-the-art methods. Besides the complete system, iKap offers a visual-to-planning network that seamlessly works with various controllers, providing a robust solution for robots navigating complex environments.
@@ I Introduction
Path planning is a fundamental task in robotics, involving the determination of a collision-free trajectory that connects the robot’s current position to the goal [1]. In dynamic and complex environments, robots must accurately perceive their surroundings and correlate environmental data with the planned path to ensure successful task execution [2]. Recently, with the development of perceptual systems, direct vision-to-planning systems have garnered increasing attention due to their potential efficiency and ability to interpret and adapt to surrounding environments [3].
@@
Data-driven end-to-end approaches, particularly those utilizing deep learning, have obtained significant advancement in recent years [3]. Such methods directly map sensor inputs to control outputs, eliminating the error propagation and latency inherent in modular systems [6, 7, 8, 3]. However, purely data-driven methods often act as black boxes, require vast amounts of data, have low sample efficiency, and exhibit a significant sim-to-real gap. In addition, these approaches lack interpretability, making it challenging to incorporate explicit physical constraints or prior knowledge [9]. Although recent works such as iPlanner [10] have explored combining data-driven methods with trajectory optimization, they neglect the robot’s kinematics. This absence of physical constraints can result in generated trajectories that are difficult for low-level controllers to execute, leading to issues such as deadlocks, collisions, or suboptimal performance [11]. As illustrated in Figure 1, for a robot with kinematic constraints like a minimum turning radius, the planned trajectory (blue curve) may not be executable by the robot.
To address the issue of end-to-end systems generating infeasible paths, we argue that kinematic constraints should be incorporated into vision-to-planning networks while retaining the end-to-end learning architecture to avoid execution errors. To this end, we propose an Imperative learning-based Kinematics-aware Planning (iKap) system that integrates the robot’s kinematic model directly into the learning pipeline. This system supervises the network to learn kinematic-feasible trajectories through self-supervised bi-level optimization (BLO). The learning module of iKap is highly flexible and can be easily integrated with various controllers to plan executable trajectories. Our main contributions include:
  • We integrate the robot’s kinematic model into the training pipeline through a differentiable model predictive control (MPC) using a self-supervised BLO framework via imperative training. This enables the network to learn the kinematically feasible trajectories.
@@ II-A Data-driven Planner
Our research builds upon iPlanner [10], an end-to-end local planner that formulates trajectory planning as a BLO problem. iPlanner guides neural networks to plan collision-free paths trained with a mixed loss with imperative learning. However, it assumes that robots can perfectly execute trajectories optimized purely from geometric loss, neglecting kinematic feasibility and thus limiting real-world performance. Additionally, iPlanner simplifies the BLO to a single-level problem via a closed-form solution for the lower-level problem, enhancing computational efficiency but reducing generalizability to other constraints. ViPlanner [12] extends iPlanner by incorporating semantic information to construct cost map for local planning, yet it does not address the limitations regarding kinematic feasibility and generalizability.
@@ II-B Differentiable Optimization and Imperative Learning
Traditional robotics algorithms often use optimization techniques to connect input and output data. Combining these optimization methods with end-to-end systems can reduce the search space for neural networks, lower the number of parameters, and boost computing efficiency [16]. Here, a differentiable optimization solver is key to ensuring proper gradient flow during learning. Amos et al. suggested implicit differentiation to handle gradient back-propagation for argmin problems, embedding it as a layer in neural networks. Later, they introduced differentiable MPC, which helps link MPC with deep learning frameworks [17, 18].
Beyond efficiency, incorporating differentiable optimization into end-to-end training frameworks also preserves the interpretability of classical methods. Building on this idea, imperative learning has been proposed, embedding optimization-based algorithms as constraints and designing the upper-level loss based on the robot’s state for self-supervised training [19]. As a result, it formulates the network training problem as a BLO. However, a key challenge lies in gradient backpropagation from the upper-level to the lower-level optimization, which many prior works have bypassed using closed-form solutions [10], gradient assumptions at convergence [20, 21], or discrete approximations [22, 23]. In our case, accurately computing gradients is critical due to the necessity of enforcing physical constraints in the BLO and the strong coupling between the two optimization levels. To address this, we employ a differentiable MPC [24, 25], demonstrating that imperative learning can effectively operate in constrained optimization scenarios.
@@ III-A Problem Formulation and System Overview
Figure 2: The training pipeline for the planner is based on BLO. In this framework, the networks, the optimized object, generate waypoints based on the given depth perception and goals. Then, the low-level trajectory optimization module tracks a kinematics-feasible trajectory. The embedded kinematics and gradients from the lower-level module are then utilized to supervise and train the networks.
@@
As illustrated in Fig. 2, our iKap consists of three key parts. First, a ResNet[26] front-end takes the depth observation $\mathcal{D}_{t}$ and encodes it into observation embedding. This embedding, combined with the target $\mathbf{p}_{t}^{g}$, is fed into the second planning network to predict a sparse key point trajectory $\boldsymbol{\mu}_{1:k}$ and collision probability $pr_{t}$. Third, a differentiable MPC tracks a reference $\boldsymbol{\tau}:=\mathbf{x}_{1:T+1}\in\mathcal{X}^{T+1}$ interpolated from $\boldsymbol{\mu}_{1:k}$ with time horizon $T$ and outputs the optimized trajectory $\boldsymbol{\tau}^{*}_{1:T+1}$ and action $\mathbf{u}^{*}_{1:T}$ while ensuring the kinematic feasibility $\mathbf{x}_{t+1}=\mathbf{F}(\mathbf{x}_{t},\mathbf{u}_{t})$. Here, $\mathbf{F}$ is the robot kinematics, $\mathbf{x}_{t}$ and $\mathbf{u}_{t}$ are the robot’s states and control input in configuration space $\mathcal{X}$ and action space $\mathcal{A}$, respectively. Finally, a well-designed traversability cost $\mathcal{U}$ is evaluated and backpropagated through both the differentiable MPC and the perception network to update the network parameters $\theta$. This process results in a BLO problem (1), where the perception and planning network is jointly optimized with the differentiable MPC.
@@
$\textrm{s.t.}\quad\mathbf{x}_{t+1}=\mathbf{F}(\mathbf{x}_{t},\mathbf{u}_{t}),$ (1c)
@@
This bi-level design separates the neural network and the traditional optimization-based approach at key-point level, allowing each to focus on its strengths—the neural network processes unstructured image inputs to predict the key-point path, while MPC generates actions and incorporates interchangeable kinematics constraints.
@@ III-C Differentiable Model-Predictive Control
$\quad\textrm{s.t.}\quad\mathbf{x}_{t+1}=\mathbf{F}(\mathbf{x}_{t},\mathbf{u}_{t}),$ (3)
@@
where $\tilde{\mathbf{x}}_{t}=\mathbf{x}_{t}-\mathbf{x}_{t}^{\text{ref}},t=0,\cdots,T$ is the error state between the current state $\mathbf{x}_{t}$ and reference trajectory $\mathbf{x}^{\text{ref}}$; $\tilde{\mathbf{x}}^{\top}_{t}\mathbf{Q}_{t}\tilde{\mathbf{x}}_{t}+\mathbf{u}_{t}^{\top}\mathbf{R}_{t}\mathbf{u}_{t}$ is the stage cost balancing the state and control effort; $\tilde{\mathbf{x}}^{\top}_{T}\mathbf{Q}_{T}\tilde{\mathbf{x}}_{T}$ is the terminal cost; $\mathbf{Q}_{t}$, $\mathbf{R}_{t}$$\mathbf{Q}_{T}$ are the weight matrices. Specifically, the reference trajectory is obtained by interpolating the waypoints predicted by the network. To show the capability of the differentiable MPC solver, we employ the classic Dubins car model (5) to represent the kinematic model of a legged robot as an example. Notice that our approach can integrate more complex kinematic models into the pipeline in a similar manner.
@@
Since $\boldsymbol{\tau}^{*}$ is a function of the reference path $\boldsymbol{\tau}$, which is a function of the predicted key point path $\boldsymbol{\mu}_{1:k}$ or $f_{\theta}$, and will be passed to the upper level optimization for the final end-to-end training, the critical part is to calculate the gradient $\frac{\partial\boldsymbol{\tau}^{*}}{\partial\boldsymbol{\tau}}$. The traditional unrolling approach maintain the computational graph throughout the entire iteration process [19], which poses significant computation burden when dealing with complex problems. It may run into divergent or vanishing. Instead, we employ the implicit function differentiation theorem and leveraging the KKT condition of (2) at the optimal point to compute the gradients of the parameters [18]. Thus, there is no need for explicit unrolling of the entire iteration process. In practice, we use the differentiable MPC module in our robot learning library PyPose [27, 24], which solves the MPC by the iterative Linear Quadratic Regulator algorithm [25]. PyPose MPC adopts a more general and problem-agnostic automatic differentiation based solution that introduces just one additional optimization iteration at the optimal point in the forward pass. This allows automatic gradient calculation during the backward pass without the need to differentiate through the entire unrolled chain or compute problem-specific analytical derivatives.
@@ III-D Upper-level Cost Design and Optimization
The cost $\mathcal{C}^{\mathcal{E}}$ in (8) is the sum of the shortest distance between the trajectory points to the environment. Each waypoint projected on a Euclidean Signed Distance Field (ESDF) map.
@@
[…]
@@
The first term encourages the planner to generate a trajectory close to the goal, and the design of $\log(\text{L2-norm}+1)$ ensures that the loss is always positive while providing a smooth gradient. The second term measures the deviation from a straight, uniformly distributed path, ensuring that the trajectory is evenly spaced. This design also helps smooth out disorganized paths, preventing the learning process from becoming trapped in local minima. The final term focuses on the relationship between the reference and kinematics-feasible trajectory. Minimizing it allows the network to learn the distribution of kinematically feasible trajectories.
To solve the BLO in (1), the gradient of the upper-level cost must be propagated back through the lower-level system by applying the chain rule, and the network parameter is updated using gradient descent:
@@ IV-A Implementation, Platforms, and Baselines
Figure 3: The real-world experiment pipeline. The experiments are conducted on the Unitree Go2 robot dog. We use AirVO to estimate the robot’s odometry, command the goal waypoint to the robot, and input this information into iKap along with depth images captured by the RealSense D435i camera.
@@
    iKap+PID: This setting is at test time. We use a PID controller to track iKap planned reference trajectory.
@@
    iKap+MPC: This setting is at test time. We use an MPC controller to track iKap planned reference trajectory.
@@ IV-B3 Navigation Test
We next evaluate the planners’ performance in a navigation task. To this end, we calculate the failure rate of a planner, where a failed attempt is defined as any instance where a collision with the environment occurred, the robot became deadlocked during planning, or the trajectory could not be executed by the controller. Specifically, we collected total 100 initial pose and goal pairs across four scenarios: "forest" (17) with dense trees for testing obstacle avoidance, "garage" (23) with partitioned warehouses, "indoor" (31) featuring narrow corridors, and "campus" (29) with diverse mixed terrains. We then used both iPlanner and iKap to generate local trajectories and navigate the robot to its destination in each environment. As shown in Table II, iKap demonstrates superior performance in three environments. It can be seen that failure rate of iKap is slightly higher than in the “Garage” environment. This is because the robot often faces obstructed vision due to the presence of numerous walls and blind spots, while iKap is sensitive to the kinematic infeasible region.
@@
iPlanner 70.59 $(\frac{12}{17})$ 73.91 $(\frac{17}{23})$ 87.10 $(\frac{27}{31})$ 79.31 $(\frac{23}{29})$
iKap 82.35 $(\frac{14}{17})$ 69.57 $(\frac{16}{23})$ 93.54 $(\frac{29}{31})$ 86.20 $(\frac{25}{29})$